OpenStudy (anonymous):

The number 1/(1 + root of 5) lies between the numbers : a) 1/3 and 1/2 b) 1/2 and 1/root2 c) 1/4 and 1/3 d) 1/5 and 1/4

OpenStudy (anonymous):

sqrt(5) lies between 2 and 3 so, 1/(1+sqrt(5)) lies between 1/3 and 1/4.

OpenStudy (anonymous):

thomas isn't there any method of finding root 5 without calculator

OpenStudy (anonymous):

help me in this one also The smallest possible integer x, for which 1260x = N^3, where N is a positive number is : a) 1470 b) 2450 c) 3675 d) 7350

OpenStudy (anonymous):

You can try some numbers: 2^2 is 4 so sqrt(5)>2 2.5^2=5.25 so sqrt(5)<2.5. You can get a bit of an idea like that.

OpenStudy (anonymous):

oh great pls help me in this one also The smallest possible integer x, for which 1260x = N^3, where N is a positive number is : a) 1470 b) 2450 c) 3675 d) 7350

OpenStudy (anonymous):

can someon help me please ?

OpenStudy (vishweshshrimali5):

Yes say kushashwa whats your problem

OpenStudy (vishweshshrimali5):

Tell me

OpenStudy (anonymous):

The smallest possible integer x, for which 1260x = N^3, where N is a positive number is : a) 1470 b) 2450 c) 3675 d) 7350 this is my problem

OpenStudy (anonymous):

see what the cube root is of 1260*1470 and 2450*1260 etc. See which is an integer.

OpenStudy (anonymous):

Does N have to be an integer?

OpenStudy (anonymous):

and one more If x is smaller then -2 , then \[\left| 1-\left| 1+x \right| \right|\]

OpenStudy (anonymous):

is it whole link

OpenStudy (anonymous):

In this case: |1+x|=|x|-1 So we get |2+|x||=|x|+2

OpenStudy (anonymous):

"is it whole link " What do you mean?

OpenStudy (anonymous):

sorry posted in wron problem but i got the solution thanks